# assumeAssumption

### v1

@@ 4.3 Matrix factorization
(i) $\Longrightarrow$ (ii) follows by Definition˜3.
(ii) $\Longrightarrow$ (iii) is due to Corollary˜2.
(iii) $\Longrightarrow$ (iv) If $\|\overline{X}\|_{2}\neq\|\overline{Y}\|_{2}$, say $\|\overline{X}\|_{2}<\|\overline{Y}\|_{2}$, then $\|t\overline{X}\|_{2}^{2}+\|t^{-1}\overline{Y}\|_{2}^{2}<\|\overline{X}\|_{2}^{2}+\|\overline{Y}\|_{2}^{2}$ for all $t\in(1,\|\overline{Y}\|_{2}^{1/2}\|\overline{X}\|_{2}^{-1/2}]$. If $(\overline{X},\overline{Y})=(0,0)$, then take $\overline{H}^{T}$ and $\overline{K}$ to be the rectangular identities. Otherwise, $\nabla^{2}f(\overline{X},\overline{Y})\neq 0$ so Theorem˜3 applies.
(iv) $\Longrightarrow$ (v) By Lemma˜5, $\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}$. Since $\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}\neq 0$, $\mathrm{Im}\hskip 0.85358pt\overline{H}^{T}=\mathrm{Im}\hskip 0.85358pt\overline{H}^{T}\overline{H}=\mathrm{Im}\hskip 0.85358pt\overline{K}\hskip 1.13809pt\overline{K}^{T}=\mathrm{Im}\hskip 0.85358pt\overline{K}\neq\{0\}$ and so $\|\overline{X}\hskip 1.13809pt\overline{Y}\|_{2}=\|\overline{X}\|_{2}\|\overline{Y}\|_{2}$ by Proposition˜9.
@@
using the same direction to obtain the equalities as in first part of the proof of Proposition˜9. Let $XY=M$. Since $f$ is $D^{2}$ and $\nabla f(X,Y)=0$, by Corollary˜2, $\accentset{\circ}{f}(X,Y,r)=\lambda_{1}(\nabla^{2}f(X,Y))r^{2}/2+o(r^{2})$. Hence, if $\lambda_{1}(\nabla^{2}f(\overline{X},\overline{Y}))<\lambda_{1}(\nabla^{2}f(X,Y))$, then there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(\overline{X},\overline{Y},r)<\accentset{\circ}{f}(X,Y,r)$ for all $r\in(0,\overline{r}]$. Otherwise, by Lemma˜5, $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}=\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}$ and so $\accentset{\circ}{f}(\overline{X},\overline{Y},r)=\accentset{\circ}{f}(X,Y,r)$ for all $r\geq 0$. In either case, $(\overline{X},\overline{Y})\preceq(X,Y)$, hence $(\overline{X},\overline{Y})$ is globally flat.
(v) $\Longleftrightarrow$ (vi) This was already shown in Lemma˜5. ∎
Proposition˜10 implies that when $(X,Y)$ is a flat global minimum, there must exist a maximal eigenvector $(H,K)$ of the Hessian such that $H^{T}H=KK^{T}$. This appears to be new. On the other hand, it is possible that $X^{T}X\neq YY^{T}$, as noted in [33]. We give an example using a well-known fact.
Fact 17.
@@
$XY=M~~~\Longleftrightarrow~~~\exists A\in\mathrm{GL}(r,\mathbb{R}):(X,Y)=(U\Sigma^{1/2}A,A^{-1}\Sigma^{1/2}V^{T}).$
@@
$XY=M$ iff $\Sigma^{-1/2}U^{T}XYV\Sigma^{-1/2}=I_{r}$ iff $\Sigma^{-1/2}U^{T}X=A$ and $YV\Sigma^{-1/2}=A^{-1}$ for some $A\in\mathrm{GL}(r,\mathbb{R})$ iff $X=U\Sigma^{1/2}A$ and $Y=A^{-1}\Sigma^{1/2}V^{T}$. ∎
Example 5.
@@
By Proposition˜10 and ˜17, $(X,Y)$ is a flat minimum iff $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$ and $(X,Y)=(M^{1/2}A,A^{-1}M^{1/2})$ for some $A\in\mathrm{GL}(2,\mathbb{R})$. Consider a decomposition $A=DNQ$ where $D=\mathrm{diag}(\alpha,\beta)$ with $\alpha,\beta>0$, $N$ is upper triangular of order 2 with ones on the diagonal with $N_{12}=\gamma\in\mathbb{R}$, and $Q\in\mathrm{O}(2)$. We have $\|X\|_{2}=\|M^{1/2}A\|_{2}=\|M^{1/2}DNQ\|_{2}=\|M^{1/2}DN\|_{2}$ and $\|Y\|_{2}=\|AM^{1/2}\|_{2}=\|Q^{-1}N^{-1}D^{-1}M^{1/2}\|_{2}=\|N^{-1}D^{-1}M^{1/2}\|_{2}$. Observe that $\|M^{1/2}DN\|_{2}=\|N^{-1}D^{-1}M^{1/2}\|_{2}=\|M^{1/2}\|_{2}=\sqrt{a}$ iff $\alpha=1$, $\sqrt{b/a}\leq\beta\leq\sqrt{a/b}$, and $N=I_{2}$. Indeed,

### v2

@@ 4.3 Matrix factorization
(i) $\Longrightarrow$ (ii) follows by Definition˜3.2.
(ii) $\Longrightarrow$ (iii) is due to Corollary˜3.19.
(iii) $\Longrightarrow$ (iv) If $\|\overline{X}\|_{2}\neq\|\overline{Y}\|_{2}$, say $\|\overline{X}\|_{2}<\|\overline{Y}\|_{2}$, then $\|t\overline{X}\|_{2}^{2}+\|t^{-1}\overline{Y}\|_{2}^{2}<\|\overline{X}\|_{2}^{2}+\|\overline{Y}\|_{2}^{2}$ for all $t\in(1,\|\overline{Y}\|_{2}^{1/2}\|\overline{X}\|_{2}^{-1/2}]$. If $(\overline{X},\overline{Y})=(0,0)$, then take $\overline{H}^{T}$ and $\overline{K}$ to be the rectangular identities. Otherwise, $\nabla^{2}f(\overline{X},\overline{Y})\neq 0$ so Theorem˜4.16 applies.
(iv) $\Longrightarrow$ (v) By Lemma˜4.23, $\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}$. Since $\overline{H}^{T}\overline{H}=\overline{K}\hskip 1.13809pt\overline{K}^{T}\neq 0$, $\mathrm{Im}\hskip 0.85358pt\overline{H}^{T}=\mathrm{Im}\hskip 0.85358pt\overline{H}^{T}\overline{H}=\mathrm{Im}\hskip 0.85358pt\overline{K}\hskip 1.13809pt\overline{K}^{T}=\mathrm{Im}\hskip 0.85358pt\overline{K}\neq\{0\}$ and so $\|\overline{X}\hskip 1.13809pt\overline{Y}\|_{2}=\|\overline{X}\|_{2}\|\overline{Y}\|_{2}$ by Proposition˜4.25.
@@
using the same direction to obtain the equalities as in first part of the proof of Proposition˜4.25. Let $XY=M$. Since $f$ is $D^{2}$ and $\nabla f(X,Y)=0$, by Corollary˜3.19, $\accentset{\circ}{f}(X,Y,r)=\lambda_{1}(\nabla^{2}f(X,Y))r^{2}/2+o(r^{2})$. Hence, if $\lambda_{1}(\nabla^{2}f(\overline{X},\overline{Y}))<\lambda_{1}(\nabla^{2}f(X,Y))$, then there exists $\overline{r}>0$ such that $\accentset{\circ}{f}(\overline{X},\overline{Y},r)<\accentset{\circ}{f}(X,Y,r)$ for all $r\in(0,\overline{r}]$. Otherwise, by Lemma˜4.23, $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}=\|\overline{X}\|_{2}=\|\overline{Y}\|_{2}$ and so $\accentset{\circ}{f}(\overline{X},\overline{Y},r)=\accentset{\circ}{f}(X,Y,r)$ for all $r\geq 0$. In either case, $(\overline{X},\overline{Y})\preceq(X,Y)$, hence $(\overline{X},\overline{Y})$ is globally flat.
(v) $\Longleftrightarrow$ (vi) $\Longleftrightarrow$ (vii) This was already shown in Lemma˜4.23. ∎
One immediately obtains the following corollary, as announced in the introduction.
Corollary 4.27.
Given $M\in\mathbb{R}^{m\times n}$, let $f:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by $f(X,Y)=\|XY-M\|_{F}^{2}/2$. If $XY=M$ and $X^{T}X=YY^{T}$, then $(X,Y)$ is flat.
Proof.
While the converse of Corollary˜4.27 is false, as noted in [38], Proposition˜4.26 implies that when $(X,Y)$ is a flat global minimum, there must exist a maximal eigenvector $(H,K)$ of the Hessian such that $H^{T}H=KK^{T}$. This appears to be new. We give a counterexample to the converse of Corollary˜4.27 using a well-known fact.
Fact 4.28.
@@
$XY=M~~~\Longleftrightarrow~~~\exists A\in\mathrm{GL}(r):(X,Y)=(U\Sigma^{1/2}A,A^{-1}\Sigma^{1/2}V^{T}).$
@@
First consider the diagonal case
$M=\begin{pmatrix}\Sigma&0\\ 0&0\end{pmatrix},~~~X=\begin{pmatrix}X_{1}\\ X_{2}\end{pmatrix},~~~\text{and}~~~Y=\begin{pmatrix}Y_{1}&Y_{2}\end{pmatrix}$.
Then $XY=M$ iff $X_{1}Y_{1}=\Sigma$, $X_{1}Y_{2}=0$, $X_{2}Y_{1}=0$, and $X_{2}Y_{2}=0$. But $X_{1}Y_{1}=\Sigma$ iff $\Sigma^{-1/2}X_{1}Y_{1}\Sigma^{-1/2}=I_{r}$ iff there is $A\in\mathrm{GL}(r)$ such that $\Sigma^{-1/2}X_{1}=A$ and $Y_{1}\Sigma^{-1/2}=A^{-1}$ iff $X_{1}=\Sigma^{1/2}A$ and $Y_{1}=A^{-1}\Sigma^{1/2}$. Now $X_{1}Y_{2}=0$ implies $Y_{2}=0$ and $X_{2}Y_{1}=0$ implies $X_{2}=0$. Thus $XY=M$ iff $X_{1}=\Sigma^{1/2}A$ and $Y_{1}=A^{-1}\Sigma^{1/2}$ and $(X_{2},Y_{2})=(0,0)$.
Now consider the general case
$M=\begin{pmatrix}U&\widetilde{U}\end{pmatrix}\begin{pmatrix}\Sigma&0\\ 0&0\end{pmatrix}\begin{pmatrix}V&\widetilde{V}\end{pmatrix}^{T}.$
Then $XY=M$ iff
$\begin{pmatrix}U&\widetilde{U}\end{pmatrix}^{T}XY\begin{pmatrix}V&\widetilde{V}\end{pmatrix}=\begin{pmatrix}\Sigma&0\\ 0&0\end{pmatrix}$
iff
$X=\begin{pmatrix}U&\widetilde{U}\end{pmatrix}\begin{pmatrix}\Sigma^{1/2}A\\ 0\end{pmatrix}=U\Sigma^{1/2}A~~~\text{and}~~~Y=\begin{pmatrix}A^{-1}\Sigma^{1/2}&0\end{pmatrix}\begin{pmatrix}V&\widetilde{V}\end{pmatrix}^{T}=A^{-1}\Sigma^{1/2}V^{T}.\qed$
Example 4.29.
@@
By Proposition˜4.26 and Fact˜4.28, $(X,Y)$ is a flat minimum iff $\|X\|_{2}=\|Y\|_{2}=\sqrt{\|M\|_{2}}$ and $(X,Y)=(M^{1/2}A,A^{-1}M^{1/2})$ for some $A\in\mathrm{GL}(2)$. Consider a decomposition $A=DNQ$ where $D=\mathrm{diag}(\alpha,\beta)$ with $\alpha,\beta>0$, $N$ is upper triangular of order 2 with ones on the diagonal with $N_{12}=\gamma\in\mathbb{R}$, and $Q\in\mathrm{O}(2)$. We have $\|X\|_{2}=\|M^{1/2}A\|_{2}=\|M^{1/2}DNQ\|_{2}=\|M^{1/2}DN\|_{2}$ and $\|Y\|_{2}=\|AM^{1/2}\|_{2}=\|Q^{-1}N^{-1}D^{-1}M^{1/2}\|_{2}=\|N^{-1}D^{-1}M^{1/2}\|_{2}$. Observe that $\|M^{1/2}DN\|_{2}=\|N^{-1}D^{-1}M^{1/2}\|_{2}=\|M^{1/2}\|_{2}=\sqrt{a}$ iff $\alpha=1$, $\sqrt{b/a}\leq\beta\leq\sqrt{a/b}$, and $N=I_{2}$. Indeed,
@@
Corollary˜4.27 means that one can compute a flat global minimum easily using linear algebra: just take a compact singular value decomposition $M=U\Sigma V^{T}$ and let $(X,Y)$ equal to $(U\Sigma^{1/2},\Sigma^{1/2}V^{T})$ up to some zero rows/columns. Below, we consider two complementary dynamical systems converging to balanced solutions, which are therefore flat. The first concerns gradient trajectories of the objective function.
Proposition 4.30.
Given $M\in\mathbb{R}^{m\times n}$ and $r\geq\mathrm{rank}\hskip 0.85358pt(M)$, let $f:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by $f(X,Y)=\|XY-M\|_{F}^{2}/2$. There exist differentiable functions $(X,Y):\mathbb{R}\to\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$ such that
$(\dot{X},\dot{Y})=-\nabla f(X,Y)$
and
$(0,0)\xleftarrow[-\infty\leftarrow t]{}(X(t),Y(t))\xrightarrow[t\to\infty]{}(X_{\infty},Y_{\infty})$
where $(X_{\infty},Y_{\infty})$ is a balanced global minimum.
Proof.
Without loss of generality, we may restrict ourselves to the diagonal case
$M=\begin{pmatrix}\Sigma&0\\ 0&0\end{pmatrix}$.
Let $(X_{0},Y_{0})=(\mathrm{diag}(\Sigma,0),\mathrm{diag}(\Sigma,0))/2\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$ and consider the initial value problem
$\left\{\begin{array}[]{ccc}(\dot{X},\dot{Y})&=&-\nabla f(X,Y),\\ (X(0),Y(0))&=&(X_{0},Y_{0}).\end{array}\right$.
It yields $\mathrm{rank}\hskip 0.85358pt(M)$ decoupled systems
$\forall i\in\llbracket 1,\mathrm{rank}\hskip 0.85358pt(M)\rrbracket,~~~\left\{\begin{array}[]{ccc}(\dot{X}_{ii},\dot{Y}_{ii})&=&-\nabla\varphi_{i}(X_{ii},Y_{ii}),\\ (X_{ii}(0),Y_{ii}(0))&=&(\Sigma_{ii}^{1/2},\Sigma_{ii}^{1/2})/2,\end{array}\right$.
where $\varphi_{i}(x,y)=(xy-\Sigma_{ii})^{2}/2$ and all other entries are fixed to zero throughout time. This reduces the problem to the scalar case, i.e., $f(x,y)=(xy-1)^{2}$, and to its dynamics on the diagonal $x=y$, i.e., on the function $g(x)=(x^{2}-1)^{2}$. We deduce that
$(0,0)\xleftarrow[-\infty\leftarrow t]{}(X_{ii}(t),Y_{ii}(t))\xrightarrow[t\to\infty]{}(\Sigma_{ii}^{1/2},\Sigma_{ii}^{1/2})$
and conclude by Corollary˜4.27. ∎
The second concerns gradient trajectories of the conserved quantity $\|C\|_{F}^{2}$. We recall a simple fact.
Fact 4.31.
Let $\phi:\mathbb{R}^{m\times n}\to\mathbb{R}$ be defined by $\phi(A)=\|A\|_{2}$. Given a singular value decomposition $A=U\Sigma V^{T}$ where $u_{i}$ and $v_{i}$ resp. denote the columns of $U$ and $V$, $\partial\phi(A)=\mathrm{co}\hskip 0.85358pt\{u_{i}v_{i}^{T}:i\in\llbracket 1,d_{A}\rrbracket\}$.
Proof.
We have $\partial\phi(A)=U\partial\phi(\Sigma)V^{T}$ and $\phi(\Sigma+tH)=\sup\{|(\Sigma+tH)w|:w\in S^{n-1}\}=\|\Sigma\|_{2}+t\max\{|H_{ii}|:i\in\llbracket 1,d_{A}\rrbracket\}=\phi(\Sigma)+t\sigma_{\{E_{ii}\}_{i\in\llbracket 1,d_{A}\rrbracket}}(H)$ where $E_{ij}$ denotes the canonical basis of $\mathbb{R}^{m\times n}$. By the max formula [43, Theorem 23.4], $\phi^{\prime}(\Sigma,H)=\sigma_{\partial\phi(\Sigma)}(H)=\sigma_{\{E_{ii}\}_{i\in\llbracket 1,d_{A}\rrbracket}}(H)$ and so $\partial\phi(\Sigma)=\mathrm{co}\hskip 0.85358pt\{E_{ii}:i\in\llbracket 1,d_{A}\rrbracket\}$. ∎
Proposition 4.32.
Given $M\in\mathbb{R}^{m\times n}$, let $f,c,\varphi:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ be defined by
$f(X,Y)$ $=\|XY-M\|_{F}^{2}/2,$
$c(X,Y)$ $=\|X^{T}X-YY^{T}\|_{F}^{2}/4,$
$\varphi(X,Y)$ $=\|X\|_{2}^{2}-2\|XY\|_{2}+\|Y\|_{2}^{2}.$
The following hold:
  1. (i)
    $f+c$ is coercive;
  2. (ii)
[…]
  3. (iii)
[…]
  4. (iv)
    For any $X_{0}Y_{0}=M$,
    $\left\{\begin{array}[]{ccc}(\dot{X},\dot{Y})&=&-\nabla c(X,Y),\\ (X(0),Y(0))&=&(X_{0},Y_{0}),\end{array}\right$.
    has a unique flattening global solution converging to a balanced global minimum of $f$ such that
    $\frac{d}{dt}\lambda_{1}(\nabla^{2}f(X,Y))\leq-(\lambda_{1}(\nabla^{2}f(X,Y))-2\|M\|_{2})^{2},$
    $\forall t>0,~~~\lambda_{1}(\nabla^{2}f(X(t),Y(t)))\leq 2\|M\|_{2}+1/t.$
Proof.
(i) Following the derivation in [24, Example 1],
$\|X\|_{2}^{4}+\|Y\|_{2}^{4}$ $\leq\|X^{T}X\|_{F}^{2}+\|YY^{T}\|_{F}^{2}$
$=\|X^{T}X-YY^{T}\|_{F}^{2}+2\langle X^{T}X,YY^{T}\rangle$
$=\|X^{T}X-YY^{T}\|_{F}^{2}+2\|XY\|_{F}^{2}$
$\leq\|X^{T}X-YY^{T}\|_{F}^{2}+2(\|XY-M\|_{F}+\|M\|_{F})^{2}$
$=2c(X,Y)+2\left(\sqrt{2f(X,Y)}+\|M\|_{F}\right)^{2}$
$\leq 2(f+c)(X,Y)+2\left(\sqrt{2(f+c)(X,Y)}+\|M\|_{F}\right)^{2}.$
(ii) For any point $(X,Y)$, we have
$\nabla c(X,Y)=\begin{pmatrix}X(X^{T}X-YY^{T})\\ -(X^{T}X-YY^{T})Y\end{pmatrix}$. (5)
If $XY=M$ and $\nabla c(X,Y)=0$, then $XX^{T}X=XYY^{T}=MY^{T}$ and $YY^{T}Y=X^{T}XY=X^{T}M$. Thus $X^{T}XX^{T}X=X^{T}MY^{T}=Y^{T}YY^{T}Y$ and $X^{T}X=YY^{T}$ by Fact˜4.18.
(iii) By Fact˜4.31, $\partial\varphi(X,Y)$ is the convex hull of the points of the form $2(s_{x}u_{x}v_{x}^{T},s_{y}u_{y}v_{y}^{T})$ where $s_{x}=\|X\|_{2}$ and $s_{y}=\|Y\|_{2}$ and $u_{x},v_{x},u_{y},v_{y}$ are some singular vectors of $X$ and $Y$ resp. The max formula [43, Theorem 23.4] states that $\varphi^{\prime}(X,Y;D)=\max\{\langle H,D\rangle:H\in\partial\varphi(X,Y)\}$. We thus compute
$~\langle(s_{x}u_{x}v_{x}^{T},s_{y}u_{y}v_{y}^{T}),\nabla c(X,Y)\rangle$
$=$ $~\langle s_{x}u_{x}v_{x}^{T},X(X^{T}X-YY^{T})\rangle-\langle(X^{T}X-YY^{T})Y,s_{y}u_{y}v_{y}^{T}\rangle$
$=$ $~s_{x}u_{x}^{T}X(X^{T}X-YY^{T})v_{x}-s_{y}u_{y}^{T}(X^{T}X-YY^{T})Yv_{y}$
$=$ $~s_{x}^{2}v_{x}^{T}(X^{T}X-YY^{T})v_{x}-s_{y}^{2}u_{y}^{T}(X^{T}X-YY^{T})u_{y}$
$=$ $~s_{x}^{4}-s_{x}^{2}|Y^{T}v_{x}|^{2}-s_{y}^{2}|Xu_{y}|^{2}+s_{y}^{4}$
$=$ $~s_{x}^{4}-|Y^{T}X^{T}u_{x}|^{2}-|XYv_{y}|^{2}+s_{y}^{4}$
$\geq$ $~s_{x}^{4}-2s_{xy}^{2}+s_{y}^{4}$
$\geq$ $~(s_{x}^{2}-2s_{xy}+s_{y}^{2})^{2}/2$
where the last inequality is due to $s_{xy}=\|XY\|_{2}\leq\|X\|_{2}\|Y\|_{2}=s_{x}s_{y}$. Indeed,
$2s_{x}^{4}-4s_{xy}^{2}+2s_{y}^{4}-(s_{x}^{2}-2s_{xy}+s_{y}^{2})^{2}$ $=s_{x}^{4}+s_{y}^{4}-8s_{xy}^{2}+4s_{x}^{2}s_{xy}+4s_{xy}s_{y}^{2}-2s_{x}^{2}s_{y}^{2}$
$=(s_{x}^{2}-s_{y}^{2})^{2}+4s_{xy}(s_{x}^{2}+s_{y}^{2}-2s_{xy})$
$\geq(s_{x}^{2}-s_{y}^{2})^{2}+4s_{xy}(s_{x}^{2}+s_{y}^{2}-2s_{x}s_{y})$
$=(s_{x}^{2}-s_{y}^{2})^{2}+4s_{xy}(s_{x}-s_{y})^{2}.$
(iv) Since $XY$ and $f$ are conserved along trajectories of $\nabla c$ by Proposition˜4.4, $f+c$ decreases along them. By (i), they are bounded and hence globally defined according to [24, Proposition 2]. They must converge to balanced points by virtue of (ii). Let $\mu=\varphi(X,Y)=\lambda_{1}(\nabla^{2}f(X,Y))-2\|M\|_{2}$, where the second equality is due to Proposition˜4.25. Since $\varphi$ agrees with a convex function along any fixed trajectory, by [10, Proposition 2] the chain rule holds almost surely:
$\dot{\mu}=\sup_{H\in\partial\varphi(X,Y)}\langle H,(\dot{X},\dot{Y})\rangle=\sup_{H\in\partial\varphi(X,Y)}\langle H,-\nabla c(X,Y)\rangle=\varphi^{\prime}(X,Y;H)\leq-\mu^{2}$
where the last inequality is due to (iii). Integrating $\dot{\mu}\mu^{-2}\geq 1$ yields $1/\mu(t)\geq 1/\mu(t)-1/\mu(0)\geq t$ if $\mu(0)\neq 0$, otherwise $\mu(t)=0$ for all $t\geq 0$. ∎
Proposition˜4.32 complements the analysis of Riemannian gradient dynamics in [21, Theorem 5.1]. Given $X_{0}Y_{0}=M$ with $\mathrm{rank}\hskip 0.85358pt(X_{0})=\mathrm{rank}\hskip 0.85358pt(Y_{0})=r$, it considers
$\left\{\begin{array}[]{ccc}(\dot{X},\dot{Y})&=&-\mathrm{grad}\hskip 0.85358pt\Phi,\\ (X(0),Y(0))&=&(X_{0},Y_{0}),\end{array}\right$.
where $\Phi:\mathrm{GL}(r)(X_{0},Y_{0})\to\mathbb{R}$ is defined by $\Phi(X,Y)=\|X\|_{F}^{2}+\|Y\|_{F}^{2}$, and the orbit is equipped with the normal Riemannian metric [21, Equation (5.3) p. 186]. [21, Theorem 5.1 (a)] shows that $\mathrm{grad}\hskip 0.85358pt\Phi=(X(X^{T}X-YY^{T}),-(X^{T}X-YY^{T})Y)$, but no connection is made with the conserved quantity. According to (5), we in fact have $\mathrm{grad}\hskip 0.85358pt\Phi=\nabla c$, which considerably simplifies the analysis and does not require the rank condition. If that condition holds, then each trajectory of $\nabla c$ remains in the orbit of the initial point.